# spacing=nonfrench

### v1

@@ IV-D Provisioning with Splitwise
Optimization. We can use three optimization goals: throughput, cost, and power. Throughput optimization is important for both the cloud service provider (CSP) and the user. The cost optimization has different importance levels to the CSP and the user. For the CSP, a higher cost for the same throughput might be acceptable, if there are gains in power and space for the datacenter. However, for the end-user, a higher cost at the same throughput is generally unacceptable. Finally, the power optimization is attractive for a CSP, since it can allow more GPU deployments in the same datacenter [36], but may not be as important to the user. We only consider the provisioned power for a machine, and not its run time power utilization in this study.
V Evaluation
V-A Methodology
Experimental setup. To evaluate our proposal on real hardware, we build Splitwise for the open-sourced vLLM [27].
Our implementation of the Splitwise technique assigns machines either a Prompt role, or a Token role. When the prompt machine generates the first token, it transfers the KV-cache to the token machine using the technique described in Section IV-C. We use MSCCL++ [9], an optimized GPU-driven communication library, to implement the naive and layer-wise KV cache transfers. Our implementation considers the contiguity of KV blocks in vLLM to minimize the number of transfers.
We run this modified vLLM on two DGX-A100s and two DGX-H100s full-machine VMs on a public cloud provider with the specifications from Table I. These are the VMs used to collect the characterization data in Section III. These machines are connected with InfiniBand and the DGX-H100s has double the bandwidth (i.e., 400 Gbps).
Simulator setup. While we use the experimental setup to evaluate the Splitwise technique, for the evaluation of the cluster-level design, we would need tens of machines. Instead, we build a simulator to explore the cluster design and evaluate our techniques. The simulator is event-driven and faithfully models the Splitwise machine pools, schedulers, machines-level memory and queues, and KV-cache transfer. For the performance model, we leverage extensive data from experiments on the DGX machines with both models across batches and request sizes. We also feed the performance impact from different prompt and generation phase mixes in a batch.
We use the prompt and token size distributions from the production traces used in Section III. As we increase and decrease the load (requests per second) for cluster sizing, we tune the Poisson arrival rate.
SLOs. To determine the maximum throughput that can be supported by a given cluster design, we use P50, P90, and P99 SLOs for TTFT, TBT, and E2E latency metrics. Table VI shows our SLO definition using DGX-A100 as a reference. We require all nine SLOs to be met. SLOs on TTFT are slightly looser, as it has a much smaller impact on the E2E latency.
@@ V-A Methodology
P50 P90 P99 TTFT 2$\times$ 3$\times$ 6$\times$ TBT 1.25$\times$ 1.5$\times$ 5$\times$ E2E 1.25$\times$ 1.5$\times$ 5$\times$
[Uncaptioned image] Figure 13: Overhead of the KV-cache transfer as the prompt size increases on A100s and H100s machines.
[Uncaptioned image] Figure 14: Overhead of KV cache transfer on TTFT, E2E latency for coding trace for H100.
Baselines. We compare our Splitwise designs against Baseline-A100 and Baseline-H100. The clusters in these baselines consist of just DGX-A100s and DGX-H100s respectively. Both baselines use the same optimized mixed continuous batching that Splitwise uses for mixed pool machines (described in Section IV-A).
V-B Experimental results
KV-cache transfer latency. We first measure the latency to transfer the KV-cache as the prompt size grows. Figure 13 shows the visible transfer latency on both A100 and H100 setups with the naive and optimized transfer design as discussed in Figure 11. Compared to the prompt computation time, the overhead is minimal ($<7\%$). The time for serialized transfers linearly increases with the prompt size since the size of the KV-cache also increases. The optimized per-layer transfer, on the other hand, hides much of the latency. For these transfers, we see a constant non-overlapped transfer time of around 8ms for the A100 and around 5ms for the H100 setup. The H100 setup has double the bandwidth of the A100 setup (i.e., 200 vs 400 Gbps), and the impact of this can be clearly seen with transfers in the H100 setup happening about twice as fast as those in the A100 setup.
As discussed in Section IV-C, for small prompt sizes ($<512$ in H100), Splitwise uses the serialized KV-cache transfer and for larger prompts, it uses per-layer transfers.
@@ V-B Experimental results
V-C Iso-power throughput-optimized clusters
@@
Cluster provisioning. We first provision the clusters using the methodology described in Section IV-D. We target a specific workload (e.g., conversation) at a peak load with the same power (i.e., iso-power) for each cluster design. For the baseline, we use the power for 40 H100 machines as our target peak power ($40\times 700W=28kW$). For the A100 baseline, we require 70 machines ($70\times 400W=28kW$). We denote these two designs as 40P/T and 70P/T as they both use mixed batching in all machines.
@@
For the Splitwise cluster designs under the coding workload trace, Splitwise-AA provisions 55 prompt machines and 15 for the token pool; denoted as (55P, 15T). Note that similarly to Baseline-A100, Splitwise-AA provisions 75% more machines than Baseline-H100. The legends in Figure 15 show the different provisioning choices under coding and conversation workloads. The design choices for longer token generation is depicted in the machine pool-sizing. For example, we provision more prompt machines for Splitwise-HH (35P, 5T) for the coding trace, while we provision more token machines (25P, 15T) for the conversation trace.
@@
Latency and throughput. Figure 15 shows a deep dive into all the latency metrics at different input load for each cluster design with the same power (i.e., iso-power). For the coding trace (Figure 14(a)), Splitwise-HH, Splitwise-HHcap, and Splitwise-AA all perform better than Baseline-H100. As the load increases, Baseline-H100 suffers from high TBT due to mixed batching with large prompt sizes. Although Splitwise-AA can support higher throughput, its TTFT is consistently higher than most designs. Splitwise-HA clearly bridges the gap by providing low TTFT and E2E at high throughput. The mixed machine pool in Splitwise becomes useful at higher loads to use all the available hardware without fragmentation. This can be clearly seen in the P50 TBT chart for Splitwise-HA, where after 90 RPS, H100 machines jump into the mixed machine pool and help run faster TBT.
@@
For the conversation trace (Figure 14(b)), Splitwise-HHcap clearly does better on all fronts, including latency. This is because the token generation phase runs for much longer than in the coding trace which is beneficial for the token machines.
@@
Figure 15: Latency metrics across input loads for iso-power throughput optimized cluster design. The dashed red lines indicate the SLO.
@@
Impact on batched tokens. Figure 16 shows the cumulative distribution of the time spent upto a given number of batched active tokens for a iso-power throughput-optimized cluster running the conversation trace at a low (70 RPS) and a high (130 RPS) load.
@@
At low load, all the 40 Baseline-H100 machines spend 70% of the time running $\leq 15$ tokens, and the rest running mixed batches with large prompts, affecting TBT and E2E. The 35 Splitwise-HH prompt machines are mostly idle, and when active, run much larger batches of tokens. The 15 Splitwise-HH token machines also do a better job at batching. Overall, Splitwise machines have a higher batching and better latency at 70 RPS. At high load, as the mixed pool comes into play, the batch sizes start looking similar across the prompt and token machines.
@@
Summary plot. Figure 16(a) summarizes the results across all cluster metrics for iso-power throughput-optimized designs for the conversation trace. We use Baseline-A100 as baseline. Compared to Baseline-A100, Splitwise-AA delivers 2.15$\times$ more throughput at the same power and cost. Compared to Baseline-H100, Splitwise-HA delivers 1.18$\times$ more throughput at 10% lower cost, and the same power.
@@
V-D Other cluster optimizations
@@
Iso-cost throughput-optimized. Figure 16(b) shows the summary plot for iso-cost clusters, with their space, throughput, and power requirements. We find that Splitwise-AA gives the best throughput for the same cost, namely 1.4$\times$ more throughput than Baseline-H100, running at 25% more power, and 2$\times$ the space. This is an interesting data point, since most customers may not care about power and space, and prefer the 40% extra throughput using older, more easily available GPUs. Whereas the answer for the CSP is not as clear.
@@
Iso-throughput power-optimized. Figure 17(a) shows cluster designs yielding same throughput at the least power. Splitwise-HHcap can achieve the same throughput as Baseline-H100 at 25% lower power at the same cost and space. This can be a clear win for the CSPs.
@@
Iso-throughput cost-optimized. Figure 17(b) shows the cost-optimized versions of the iso-throughput design. Note that there are no changes to any of the homogeneous designs between Figure 17(a) and Figure 17(b). This is because the prompt and token machines have the same cost and power. Splitwise-HA and Splitwise-HHcap though, arrive at slightly different results with the cost and power optimizations. Figure 17(b) shows that with Splitwise-AA, customers can achieve the same throughput as Baseline-H100, at 25% lower cost.
Figure 19: Latency impact of running a workload on a cluster designed for another workload. The dashed red lines indicate the SLO.
V-E Impact of workload changes
@@
So far, we have tested a trace and a model on clusters optimized for a specific workload pattern and model. To test the robustness of Splitwise, we now run conversation trace on a cluster meant for coding service, and Llama2-70B on a cluster meant for BLOOM-176B. Figure 19 shows these results for iso-power throughput-optimized clusters.
@@
Changing workload trace. Compared to Figure 14(b), we find that in Figure 18(a), the Baseline clusters are similarly sized and see no throughput or latency impact. Splitwise-AA and Splitwise-HH with the mixed pool based scheduling morph well to meet the requirements of the new workload, and see no throughput or latency impact. Since Splitwise-HA and Splitwise-HHcap have different types of machines for Prompt and Token, they experience a throughput setback of 7% from the respective cluster optimized designs for conversation trace. Note that all the Splitwise designs still perform much better than any of the Baseline designs.
@@
Changing model. Figure 18(b) shows that Llama2-70B can support much higher throughput in the same cluster design than BLOOM-176B, given its fewer parameters (Table III). All the Splitwise designs out-perform both the Baseline designs at higher load. Furthermore, Splitwise-HH and Splitwise-HHcap consistently achieve the best latency, even as the load increases.
@@
Summary. Based on these two experiments, we conclude that Splitwise can morph according to the requirements of the workload using its smart scheduling, and is robust to changes in the LLM models, load, and token distributions.
V-F Cluster design for batch job
@@
We design various clusters with Splitwise under strict latency SLOs, even when we are optimizing for throughput. This might not be necessary for batch jobs, that can be stressed to high load, while measuring the generation throughput without latency requirements. We find that upon stressing our iso-power throughput-optimized clusters, Baseline-A100 and Splitwise-AA have the best throughput per cost at 0.89 RPS/\$. At high load, Splitwise devolves into the iso-count Baseline, since it starts mixed batching with all the machines in the mixed pool. The same holds true for Splitwise-HH and Baseline-H100, that achieve 0.75 RPS/\$.
VI Discussion
Other hardware for the token phase. In this work, the only hardware variants we use are A100 and H100. However, given the token phase characteristics, GPUs like AMD MI-250 [1] and CPUs like Intel Sapphire-Rapids (with HBM) [7] could be effective token machines due to their high memory capacities and bandwidths.
What would optimized hardware look like? Based on our characterization, the prompt phase would need high compute capability with high enough memory bandwidth, but not necessarily a lot of memory capacity. The token phase would need high-enough compute capacity with high memory capacity and bandwidth. Splitwise enables this hardware design space exploration for each phase independently.
@@ VI Discussion
Interconnect between prompt and token machines. In this work, we assume Infiniband connection between the prompt and token machines in all the designs (albeit, lower bandwidth when A100s were involved). Although this is common in almost all cloud environments, one of the designs, Splitwise-HA might not be readily available with an Infiniband connection between H100s and A100s machine pools, even though technically feasible. The alternative could be HPC clouds, with Infiniband connections through the CPU [2], or ethernet, using RoCE [33]. Given our optimized KV-cache transfer in smaller blocks utilizing less bandwidth and reducing critical latency however, an interconnect with 10×10$\times$10 × less bandwidth would still be beneficial. To further reduce our bandwidth utilization, we could also compress the KV-cache before transferring it across the network [30].
@@
Heterogeneous prompt/token machines. Although Figure 19 shows that the mixed-pool based scheduling of Splitwise makes it amenable to varied models and input traces, we recognize that fragmenting a datacenter with different types of GPUs may bring its own challenges for the CSP.
@@
VII Related Work
@@
LLM serving. LLM inference serving is a fast developing area, with several recent works optimizing the batching, caching and scheduling [48, 23, 20, 28, 27, 19, 39, 18, 47, 44]. Several prior works have also proposed using CPUs and devices with lower compute capability [6, 10]. However, all the prior work uses the same machine for prompt and token phases. With Splitwise, these systems could benefit by increasing throughput due to phase splitting.
Heterogeneous scheduling. Much prior work exists in the field of heterogeneous scheduling for interactive services [49, 38, 26]. These efforts typically exploit hardware heterogeneity to strike a desirable balance across multiple user and/or CSP objectives such as cost, energy, and performance. Splitwise currently leverages mixed pools for better scheduling, and we plan to explore other scheduling techniques in our future work.
VIII Conclusion
@@
In this work, we extensively characterized the prompt and token generation phases of LLM inference and highlighted the difference in their system utilization patterns. Based on insights from the characterization, we designed Splitwise, which enables separate machines to run the two phases and increases utilization. It opens up a new exploration space as the machine pools for the two phases can be designed and scaled separately. We explored Splitwise designs optimized for throughput, cost, and power, and demonstrated that it performs well even as workloads change. We show that with Splitwise under latency SLOs, we can achieve 1.76×1.76$\times$1.76 × better throughput with 15% lower power at the same cost, or 2.35×2.35$\times$2.35 × better throughput with same cost and power.

### v2

@@ V-B Simulator setup
[…]
@@
Baselines. We compare our Splitwise designs against Baseline-A100 and Baseline-H100. The clusters in these baselines consist of just DGX-A100s and DGX-H100s, respectively. Both baselines use the same mixed continuous batching that Splitwise uses for mixed pool machines (described in Section IV-A).
VI Evaluation
VI-A Experimental results
[…]
As discussed in Section IV-C, for small prompt sizes (<512absent512<512< 512 in H100), Splitwise uses the serialized KV-cache transfer and for larger prompts, it uses per-layer transfers.
Figure 14: Overhead of the KV-cache transfer as the prompt size increases on A100s and H100s.
@@ VI-A Experimental results
Figure 15: Overhead of KV cache transfer on TTFT, E2E latency for coding trace for A100 and H100.
VI-B Iso-power throughput-optimized clusters
@@
Cluster provisioning. We provision clusters using the methodology described in Section IV-D. We target a specific workload (e.g., conversation) at a peak load with the same power (i.e., iso-power) for each cluster design. For the baseline, we use the power for 40 DGX-H100 machines as our target peak power. For the A100 baseline, we can fit 70 DGX-A100 machines under the same power budget. We denote these two designs as 40P/T and 70P/T respectively, since they both use mixed batching in all machines.
@@
For Splitwise cluster designs under the coding trace, Splitwise-AA provisions 55 prompt machines and 15 for the token pool, denoted as (55P, 15P). Note that like Baseline-A100, Splitwise-AA also provisions 75% more machines than Baseline-H100. The legends in Figure 16 show the different provisioning choices under coding and conversation workloads. Request size distributions reflect in the machine pool sizing. For example, we provision more prompt machines under Splitwise-HH (35P, 5T) for the coding trace, while we provision more token machines (25P, 15T) for the conversation trace.
@@
Latency and throughput. Figure 16 shows a deep dive into all the latency metrics at different input load for each cluster design with the same power (i.e., iso-power). For the coding trace (Figure 16(a)), Splitwise-HH, Splitwise-HHcap, and Splitwise-AA all perform better than Baseline-H100. As the load increases, Baseline-H100 suffers from high TBT due to mixed batching with large prompt sizes. Although Splitwise-AA can support higher throughput, its TTFT is consistently higher than most designs. Splitwise-HA clearly bridges the gap by providing low TTFT and E2E at high throughput. The mixed machine pool in Splitwise becomes useful at higher loads to use all the available hardware without fragmentation. This benefit can be seen clearly in the P50 TBT chart for Splitwise-HA, where after 90 RPS, H100 machines jump into the mixed machine pool and help reduce TBT.
@@
For the conversation trace (Figure 16(b)), Splitwise-HHcap clearly does better on all fronts, including latency. This is because its token generation phases typically run for much longer than in the coding trace, which is beneficial for the token machines.
@@
Figure 16: Latency metrics across input loads for iso-power throughput optimized clusters. Dashed red lines indicate SLO.
@@
Impact on batched tokens. Figure 17 shows the cumulative distribution of time spent processing a varying number of batched active tokens in an iso-power throughput-optimized cluster. The distributions are collected by running the conversation trace at low (70 RPS) and high (130 RPS) loads.
@@
At low load, all 40 Baseline-H100 machines spend 70% of the time running $\leq$15 tokens, and the rest running mixed batches with large prompts, which affects TBT and E2E. The 35 Splitwise-HH prompt machines are mostly idle, and when active, run much larger batches of tokens. The 15 Splitwise-HH token machines also do a better job at batching. Overall, Splitwise machines have better batching and latency at 70 RPS. At high load, since the mixed pool is utilized more, the batch sizes start looking similar across prompt and token machines.
@@
[…]
@@
VI-C Other cluster optimizations
@@
[…]
@@
Iso-throughput power-optimized. Figure 19(a) shows cluster designs that yield same throughput at the least power. Splitwise-HHcap can achieve the same throughput as Baseline-H100 at 25% lower power at the same cost and space. This can be a clear win for the CSPs.
@@
Iso-throughput cost-optimized. Figure 19(b) shows the cost-optimized versions of the iso-throughput design. Note that there are no changes to any of the homogeneous designs between Figures 19(a) and 19(b). This is because the prompt and token machines have the same cost and power. However, Splitwise-HA and Splitwise-HHcap arrive at slightly different results with the cost and power optimizations. Figure 19(b) shows that with Splitwise-AA, customers can achieve the same throughput as Baseline-H100 at 25% lower cost.
VI-D Impact of workload changes
@@
So far, we have tested a trace and a model on clusters optimized for a specific workload pattern and model. To test the Splitwise’ robustness, we now run conversation trace on a cluster meant for coding service, and Llama-70B on a cluster meant for BLOOM-176B. Figure 20 shows these results for iso-power throughput-optimized clusters.
@@
Changing workload trace. Compared to Figure 16(b), we find that in Figure 20(a), the Baseline clusters are similarly sized and incur no throughput or latency impact. Splitwise-AA and Splitwise-HH with the mixed pool morph well to meet the requirements of the new workload, and they see no throughput or latency impact. Since Splitwise-HA and Splitwise-HHcap have different types of machines in the prompt and token pools, they experience a throughput setback of 7% from the respective cluster optimized designs for conversation trace. Note that all the Splitwise designs still perform much better than any of the Baseline designs.
@@
Changing model. Figure 20(b) shows that Llama-70B can support much higher throughput in the same cluster design than BLOOM-176B, given its fewer parameters (Table III). All the Splitwise designs out-perform both the Baseline designs at higher load. Furthermore, Splitwise-HH and Splitwise-HHcap consistently achieve the best latency, even as the load increases.
@@
Summary. Based on these two experiments, we conclude that Splitwise can morph according to the requirements of the workload using its smart scheduling, and it is robust to changes in the LLMs, request load, and token distributions.
Figure 20: Latency impact of running a workload on a cluster designed for another workload. Dashed red lines indicate SLO.
VI-E Cluster design for batch job
@@
We design various clusters with Splitwise under strict latency SLOs, even when we are optimizing for throughput. This is unnecessary for batch jobs, which can be stressed to high load for a high token generation throughput. We find that upon stressing our iso-power throughput-optimized clusters, Baseline-A100 and Splitwise-AA have the best throughput per cost at 0.89 RPS/\$. At high load, Splitwise devolves into the iso-count Baseline, since it starts mixed batching with all the machines in the mixed pool. The same holds true for Splitwise-HH and Baseline-H100, which achieve 0.75 RPS/\$.
VII Discussion
Extensibility to new models. Despite the plethora of model sizes from 2B parameters  [47, 84] to 176B parameters [69] or more [18], all modern transformer-based generative LLMs have the distinct prompt processing and token generation phases. Similarly, even modifications and flavors like Mixture-of-Experts (MoEs) have these phases. Since Splitwise is built solely by exploiting these phases, it is applicable to all of the current and upcoming LLMs, as long as the auto-regressive nature of the workload requires these two phases. Note that as shown in Section VI-D, clusters provisioned with Splitwise for one model can also efficiently serve other models.
Alternative compute hardware. In this work, we use NVIDIA H100 and A100 GPUs since they are commonly used for LLM inference in datacenters today [17]. Smaller datacenter GPUs like NVIDIA T4 lack enough memory to run modern LLMs efficiently. In general, our methodology is applicable to any hardware (including CPUs, FPGAs, ASICs [33]) that aligns with the computational requirements of prompt and token phases. Our characterization suggests that prompt phases need high compute capability and memory bandwidth with low memory capacity, whereas token phases need moderate compute capability with high memory capacity and bandwidth. Thus, GPUs like AMD MI-250 [2] and CPUs like Intel Sapphire-Rapids (with HBM) [9] could be effective token machines. Since we do not have access to such hardware and/or optimized LLM implementations, we leave this to future work.
@@ VII Discussion
Interconnect between prompt and token machines. In this work, we assume Infiniband connection between the prompt and token machines in all the designs (albeit, lower bandwidth when A100s were involved). Although this is common for all homogenous machines, Splitwise-HA is not be readily available with an Infiniband connection between H100s and A100s, even though technically feasible. The alternative could be HPC clouds, with Infiniband connections through the CPU [3], or Ethernet, using RoCE [58]. Given our optimized KV-cache transfer that helps reduce critical latency, an interconnect with $10\times$10 × lower bandwidth would likely still be beneficial. To further reduce our bandwidth utilization, we could also compress the KV-cache before transferring it across the network [55].
@@
Heterogeneous prompt/token machines. Although Splitwise is robust to varied models and input traces, we recognize that fragmenting a data center with different types of GPUs (e.g., Splitwise-HA) may bring its own challenges for the CSP.
@@
VIII Related Work
Heterogeneous scheduling and dataflow systems. Prior work has studied heterogeneous scheduling for a variety of interactive services [83, 65, 68]. These works exploit hardware heterogeneity to strike a balance between different objectives such as cost, energy, and performance. However, they run the entire workload on the same machine. Research on heterogeneous multiprocessor CPU scheduling attempts to match workload heterogeneity to hardware heterogeneity [40, 76, 41, 50, 80, 29]. These works use profiling or online monitoring with metrics like request length or hardware performance counters to identify workload phases and allocate them appropriately on heterogeneous processors. However, they do not consider the complexities with batching. Distributed dataflow systems orchestrate large-scale computational graphs and aim to provide general-purpose programmability [34, 46, 75, 82]. LLM inference under Splitwise can be viewed as a static computational graph with two stages, so it could be implemented using distributed frameworks that provide efficient GPU abstractions [59]. Splitwise differs from these works since it uses a specialized two-phase design for generative LLM inference and leverages phase-aware resource management with efficient batching.
@@ VIII Related Work
Model serving systems. LLM inference serving is a rapidly developing field, with several recent works optimizing batching [81, 51, 53, 23, 25], scheduling [51, 66, 73, 79, 22, 42], and memory usage [32, 51, 35, 74]. Prior work has also proposed using CPUs and lower compute capability devices for LLM serving [8, 12]. These approaches use the same machine for both prompt and token phase. With Splitwise, they could improve throughput and latency by splitting phases.
Prior work on video and ML serving focuses on scheduling model chains with data dependencies under latency constraints [43, 31, 68, 49, 24]. Such schedulers rely on model profiling to make efficient allocation decisions and manage requests across machines. Recommendation system inference exhibits compute/memory heterogeneity both within and across models. Prior work exploits such heterogeneity to selectively schedule requests between CPUs and accelerators [38, 52], colocate models with complementary memory usage [30], and partition compute/memory on heterogeneous hardware resources [45, 48]. Similarly, Splitwise exploits the heterogeneity within LLM inference requests. However, it uses different optimizations due to the differences in LLM workload characteristics and requirements.
IX Conclusion
@@
We extensively characterized the prompt computation and token generation phases of LLM inference to draw out differences in their system utilization patterns. Based on our insights, we designed Splitwise to separate these phases onto different machines and enable phase-specific resource management. Using Splitwise, we explored cluster designs optimized for throughput, cost, and power, and showed that they perform well even as workloads change. Splitwise clusters under performance SLOs achieve $1.76\times$1.76 × better throughput with 15% lower power at the same cost, or $2.35\times$2.35 × better throughput with same the cost and power than existing designs.
Acknowledgements
We thank the reviewers for their helpful feedback. We thank Chetan Bansal, Srikant Bhardwaj, Suriya Kalivardhan, Ankur Mallick, Deepak Narayanan, and Amar Phanishayee for insightful discussions. Pratyush Patel was partially supported by NSF CNS-2104548 and a research grant from VMware.
Appendix A Artifact Appendix
A-A Abstract
We open source critical components needed to evaluate Splitwise; these could be repurposed to also evaluate future LLM inference serving systems. Our artifact includes:
  • Production traces from two LLM inference services at Microsoft Azure.
  • A prototype implementation of Splitwise’s KV-cache transfer mechanism in vLLM [51].
  • SplitwiseSim, a discrete event simulator to evaluate model serving in LLM inference clusters.
Artifact functionality was only tested for the traces and SplitwiseSim due to limited hardware availability.
A-B Artifact check-list (meta-information)
  • Data set: Production traces available as a part of the artifact.
  • Run-time environment: Linux / Ubuntu.
  • Hardware: Two machines connected over GPU Infiniband for the vLLM prototype (e.g. NVIDIA DGX-A100, NVIDIA DGX-H100). x86-64 CPU machine for SplitwiseSim.
  • Publicly available?: Yes.
  • Code licenses (if publicly available)?: MIT.
  • Data licenses (if publicly available)?: CC-BY.
  • Archived (provide DOI)?: 10.5281/zenodo.11003049.
A-C Description
How to access. The entire artifact is available as an archive on Zenodo: https://doi.org/10.5281/zenodo.11003049. Individual components are also available online as follows:
  • The production traces can be downloaded from the Azure Public Dataset GitHub repository [4].
  • The KV-cache transfer prototype can be downloaded from the vLLM GitHub repository, currently available as a pull request [1].
  • SplitwiseSim, and the associated experiment and plotting scripts, can be downloaded from a separate GitHub repository [20].
Hardware dependencies. The KV-cache transfer prototype requires two GPU machines connected over Infiniband, such as NVIDIA DGX-A100s or NVIDIA DGX-H100s. SplitwiseSim requires a standard x86-64 CPU machine; multiple machines may be used to parallelize simulation runs.
Software dependencies. The KV-cache transfer prototype is built on top of vLLM [51] and MSCCL++ [11]. SplitwiseSim depends on a small set of publicly available Python packages, which can be installed via the included requirements.txt.
Data sets. Coding and conversation traces from Microsoft Azure are available online as a part of the artifact release [4].
A-D Installation and Experiment Workflow
Please refer to the README files within the artifact for installation and usage instructions.
A-E Methodology
Submission, reviewing and badging methodology: